# A Unified Framework for N-Dimensional Visualization and Simulation: Implementation and Evaluation including 4D Boolean Operations

### v2

@@ Abstract
This study proposes a unified framework for simulation and visualization of intuitive exploration of phenomena in $N$-dimensional space. While specialized libraries offer powerful geometric algorithms, they typically lack integrated environments for interactive trial and error, creating a barrier for researchers. The contribution of this research is the integration of Quickhull-based mesh generation, visualization via hyperplane slicing, and computationally expensive Boolean operations into a single, extensible platform, while maintaining interactivity. To validate its effectiveness, this paper presents a 4-dimensional implementation and introduces a new interaction design, termed ‘High-Dimensional FPS,’ to enable intuitive high-dimensional exploration. Furthermore, as a case study to demonstrate the framework’s high extensibility, I also integrated a non-rigid body physics simulation based on Extended Position Based Dynamics (XPBD). Experimental results confirmed the effectiveness of the proposed method, achieving real-time rendering (80 fps) of complex 4D objects and completing Boolean operations within seconds in a standard PC environment. By providing an accessible and interactive platform, this work lowers the entry barrier for high-dimensional simulation research and enhances its potential for applications in education and entertainment.
@@ 1 Introduction
While the simulation and visualization of high-dimensional spaces are theoretically established to some extent, practical integrated environments and concrete methods are limited. This study proposes a unified framework for $N$-dimensional simulation and visualization that can be easily adopted and applied by researchers and developers. This framework unifies previously disparate techniques—such as Quickhull-based high-dimensional convex hull mesh generation, visualization based on hyperplane slicing, high-dimensional exploration through view transformations, and mesh processing via Boolean operations—enabling their integrated use and providing an environment for the intuitive understanding of high dimensions.
Previously, functionalities such as high-dimensional mesh generation and Boolean operations were offered within specialized computational geometry libraries like the Computational Geometry Algorithms Library (CGAL). However, an environment that combined these with visualization and physics simulation to enable interactive trial and error has been lacking.
The framework proposed in this paper is positioned as one solution to this challenge. Its architecture, founded on the design principle of separating topology and geometry, achieves high extensibility. I demonstrate its effectiveness through a case study integrating a non-rigid body physics simulation based on Extended Position Based Dynamics (XPBD), showing that external components can be easily integrated. This approach is expected to lower the barrier to entry for high-dimensional simulation research, enhance its potential for applications in education and entertainment, and contribute to the future growth of the research community and academic advancement.
@@ 2 Related Work
Prior research in high-dimensional simulation and visualization has primarily focused on reporting individual techniques. In mesh generation, the construction of convex hull meshes using the Quickhull algorithm is widely known [quickhull], but it has a high computational cost to guarantee mathematical rigor. In physics simulation, while simulations of non-rigid bodies based on XPBD have been reported for 3D space [XPBD], and research on the simulation of rigid bodies in 4D and higher dimensions exists [nrigid], there are exceedingly few reports of integrated environments that allow for the interactive manipulation of non-rigid bodies and the visualization of objects in 4D or higher dimensions. Turning to mesh editing, Boolean operations are a standard technique in 3D. In higher-dimensional spaces, specialized computational geometry libraries like CGAL also provide this functionality as an offline computation. However, these processes require significant computational resources to guarantee mathematical precision and are not intended for interactive use that involves real-time result verification. Regarding visualization, while a GPU-based 4D visualization architecture has been proposed [GL4D], its primary focus is on fast rendering, and it offers a limited environment for the kind of exploratory experience that this paper aims to achieve.
@@ 3.1 Design Philosophy and Architectural Overview
This section describes the technical details of the proposed unified framework. This framework integrates the functionalities of mesh generation, Boolean operations, visualization, and physics simulation into a single, extensibility-focused pipeline. I will first provide an overview of the entire architecture, and then detail the core algorithms that constitute each function and their implementation.
In the design and implementation of this framework, three principles were emphasized: algorithmic transparency, minimal reliance on external libraries, and future extensibility. This is a design decision that prioritizes the framework’s utility as a research platform—allowing researchers to easily understand, verify, and incorporate new functionalities into the high-dimensional geometry algorithms—over extreme performance optimization. Based on this philosophy, all computational geometry logic is unified in a standard C# implementation on the CPU. While GPU-based parallelization can deliver high performance for specific tasks, it introduces issues of code complexity and hardware dependency. The CPU-centric approach of this paper aims to enhance platform independence and algorithmic transparency. This implementation strategy is expected to lower the technical barriers for other researchers to verify this study’s algorithms or apply them to their own research. Furthermore, within the scope of this paper, numerical robustness related to floating-point arithmetic is excluded from consideration and is left for future work. The software for this research was implemented in C# on the Unity engine and is available as an open-source project [Arai_2025_UnifiedFramework].
3.2 Overview of the Overall Architecture
Figure 1: An overview of the proposed framework. The left side shows the static shape definition workflow, and the right side shows the dynamic visualization pipeline.
As illustrated in Figure 1, the overall architecture and data flow of this framework consist of multiple loosely coupled components: convex hull mesh generation, mesh editing, and visualization. The mesh generation component uses a Quickhull-based algorithm to construct an index list of facets from the input point cloud data, which defines the topology of the $N$-dimensional object. The mesh editing component dynamically processes and modifies the mesh’s topology and geometry based on user operations, such as Boolean operations. The visualization component generates and renders a 3D cross-section each frame by performing hyperplane slicing on the coordinate-transformed mesh, based on the current vertex coordinates and the user’s viewpoint.
Each component interacts only through a well-defined data structure for the object’s topology and geometry, thereby ensuring high modularity. Furthermore, I have adopted Geometric Algebra as a unified mathematical foundation that enables the consistent handling of geometric entities such as $N$-dimensional vectors, planes, and rotations. For Geometric Algebra itself, which I use as my foundation, [vince2008] is cited as an excellent introductory text for researchers in computer graphics.
3.3 N-Dimensional Convex Hull Construction
@@ 3.3.1 Comparison with Existing Methods
This component is centered around a self-contained C# implementation of the classic $N$-dimensional Quickhull algorithm [quickhull]. A standard implementation would perform an optimization called candidate point pruning to maximize computational efficiency. This strategy, in which each facet individually manages a list of candidate points and excludes points absorbed into the convex hull’s interior from computation, is extremely effective for volumetric point clouds containing numerous interior points. However, the primary applications within my framework—such as the basic convex hull mesh construction for initial shapes and dynamic simplicial tessellation (detailed in Section 3.3.3)—involve frequently handling small-scale, surface-like point clouds, typically with fewer than several dozen vertices and almost no interior points. In such contexts, the performance gain offered by a pruning mechanism is limited, while the added implementation complexity reduces code maintainability and introduces risks to robustness in a floating-point arithmetic environment.
Therefore, my implementation considers this trade-off and intentionally omits the pruning mechanism. As shown in Figure 2, my algorithm performs a linear search over a single, global set of unprocessed points to find the farthest point, and removes only those points that are used as vertices of the convex hull from the set. This simplified state-management approach, in exchange for theoretical computational efficiency on large volumetric point clouds, provides the practically significant advantages of implementation simplicity and independence from external libraries for my primary use cases. Hereafter in this paper, I will refer to this method as ”Direct Quickhull.”
@@ 3.3.2 Calculating N-Dimensional Normals and Visible Facets
In Direct Quickhull, efficiently computing facet normal vectors plays a crucial role in determining overall performance. The conventional approach based on null-space computation is unsuitable for real-time applications, as it typically requires methods such as Singular Value Decomposition (SVD). To solve this problem, my implementation introduces the concepts of the wedge product and the dual from Geometric Algebra. First, in an $N$-dimensional space $\mathbb{R}^{N}$, I define a facet from its NNvertices $\{\mathbf{p}_{0},\mathbf{p}_{1},\dots,\mathbf{p}_{N-1}\}$and construct the (N1)(N-1) linearly independent vectors that span the facet.
@@
Next, I calculate the wedge product of these vectors.
@@
[…]
@@
Next, it is necessary to correctly determine the orientation of the facet’s normal vector. In conventional 3D, the correct orientation of a normal vector for a facet on the surface of a convex hull was obtained by computing the vertex winding order, which relies on the right-handed rule of the cross product. However, the cross product is unique to 3D space. In higher dimensions, there are no geometric concepts analogous to the winding order that intuitively define a face’s orientation.
To solve this problem, my implementation uses the geometric property that the normal vectors of a convex hull’s facets must always point outward. This orientation is determined by calculating the dot product of the facet’s normal vector and a vector pointing from a point on the facet to the centroid of the initial simplex constructed at the beginning of the algorithm. If the sign of the dot product is negative, the normal vector is determined to be pointing outward. If the sign is positive, its sign is inverted, thereby ensuring that I always obtain a normal vector with the correct outward-facing orientation.
@@
Next, I use the facet normal vectors to determine the visible facets, which is a necessary step in the process of expanding the convex hull.
@@
3.3.3 Method for Simplicial Tessellation during Convex Hull Construction
my implementation not only computes the facets that constitute the surface of the convex hull but is also equipped with the functionality to simultaneously generate its interior volume as a set of $N$-dimensional simplices. This is accomplished by applying the incremental process of the Quickhull algorithm. In Quickhull’s incremental process, the newly added volume is represented as a set of $N$-dimensional simplices formed from the new vertex $\mathbf{p}_{\text{new}}$ and the visible facets. my implementation directly uses this principle, appending the corresponding simplices to a list. By repeating this process, I simultaneously obtain both the surface facets of the convex hull and a list of simplices that perfectly tessellates its interior volume. This simplicial decomposition data serves two important roles within this framework. Its primary application is to serve as the base data for rapidly performing a simplicial decomposition of the interior of facets clipped during Boolean operations. Furthermore, it is intended for future application in efficiently determining the cross-sectional structure of facets from hyperplane slicing in spaces of five or more dimensions, a topic detailed in Section 3.5.3.
@@ 3.4 N-Dimensional Boolean Operations
3.4.1 Algorithm Overview and Design Philosophy
While Boolean operations are an extremely powerful method for interactively modifying the shape of an $N$-dimensional mesh, their interactive implementation in high-dimensional space presents a significant challenge from a computational cost perspective. To address this challenge, my framework implements a practical four-stage algorithm. First, in a broad phase, I use spatial partitioning to narrow down potential intersection candidates. Next, in a narrow phase, I determine the actual geometric intersections. The intersected facets are then decomposed into smaller facets in a tessellation stage. Finally, the results are classified in an inside-outside test stage to construct the final operation result.
In designing this algorithm, I prioritized a balance of three factors over extreme performance optimization or a full guarantee of mathematical rigor: implementation simplicity, extensibility to $N$-dimensions, and a practical level of performance that enables interactive trial and error. The following sections will detail the specific techniques used in each stage and the design decisions behind them.
@@ 3.4.2 Broad Phase: Candidate Pruning via Spatial Partitioning
In $N$-dimensional Boolean operations, a brute-force intersection test between all facets of two meshes, with $N_{\text{A}}$ and $N_{\text{B}}$ facets respectively, has a computational complexity of $O(N_{\text{A}}N_{\text{B}})$, making it a major bottleneck in interactive implementations. To avoid this problem, my implementation introduces a spatial partitioning grid as the broad phase to efficiently prune pairs of facets that could potentially intersect. I compute the axis-aligned bounding box (AABB) for each facet and register the facet with the grid cells that its AABB occupies. As a result, only facet pairs that share a common cell become candidates for the precise intersection tests in the subsequent narrow phase. Compared to hierarchical spatial partitioning methods such as BVH or Octree, the uniform grid approach adopted in my implementation offers the advantage of being simple to implement while still achieving practical pruning performance for my target use cases.
@@ 3.4.3 Narrow Phase: Facet-Facet Intersection Test
In the narrow phase, I perform precise geometric intersection tests on the candidate pairs pruned during the broad phase. The principle of this algorithm is a generalization to $N$-dimensions of the method proposed by Tomas Möller et al. for fast triangle-triangle intersection tests in 3D [Moller1997]. Its essence lies in decomposing the intersection test between two facets, $\mathcal{F}_{\text{A}}$ and $\mathcal{F}_{\text{B}}$, into two symmetric sub-problems. First, I compute the geometric intersection region, $\mathcal{I}_{\text{B}}$, between the (N1)(N-1)-dimensional hyperplane spanned by $\mathcal{F}_{\text{A}}$ and the facet $\mathcal{F}_{\text{B}}$. Second, I perform the same operation with the roles swapped, computing the intersection region $\mathcal{I}_{\text{A}}$ between the hyperplane spanned by $\mathcal{I}_{\text{B}}$ and the facet $\mathcal{F}_{\text{A}}$. Both of these intersection regions are (N2)(N-2)-dimensional convex polytopes. Finally, I determine if an intersection between the facets exists by testing whether $\mathcal{I}_{\text{A}}$ and $\mathcal{I}_{\text{B}}$ overlap within their common (N2)(N-2)-dimensional space.
If an intersection is found to exist, this information is utilized in the subsequent tessellation stage. Specifically, I register this intersection pair in a dictionary where one facet is the key and the other is the value. By performing this process for all intersection pairs, I can systematically construct information about the clipping relationships, enabling us to handle complex situations where one facet intersects with multiple other facets.
In my 4D implementation, the intersection region is a 2D convex polygon. Since the number of vertices is small, I adopted the Separating Axis Theorem (SAT), which is relatively straightforward to implement. However, in five or more dimensions, the maximum number of vertices of the (N2)(N-2)-dimensional polytope intersection region increases, making the implementation of SAT more complex. Therefore, for extension to higher dimensions, an intersection test approach using the GJK algorithm is considered a more practical option.

### v3

# A Unified Architecture for N-Dimensional Visualization and Simulation: 4D Implementation and Evaluation including Boolean Operations

@@ Abstract
This paper proposes a unified software architecture for visualization and simulation based on a design targeting an $N$-dimensional space. The contribution of this work lies in presenting an architectural configuration that integrates multiple processes into a single software architecture: Quickhull-based convex hull mesh generation, Boolean operations, coordinate transformations for high-dimensional exploration (including orientation and view transformations), and hyperplane slicing for visualization. The proposed approach adopts an approximate implementation that tolerates numerical errors and prioritizes implementation transparency over guarantees of numerical rigor. The experimental results and evaluations presented in this paper are limited to a 4D implementation; no evaluation is conducted for $N>4$, and the discussion is restricted to stating that the architecture itself has a dimension-independent structure. This paper also proposes an interaction design for high-dimensional exploration based on FPS navigation. As an input example involving shape changes over time, a non-rigid body simulation based on XPBD (Extended Position Based Dynamics) is integrated into the 4D implementation. Experimental results confirm that the 4D implementation runs on a single PC.
@@ 1 Introduction
This paper builds on prior work on visualization and simulation in high-dimensional spaces. However, examples remain limited that integrate Quickhull-based convex hull mesh generation, Boolean operations, coordinate transformations for high-dimensional exploration (orientation and view transformations), and hyperplane slicing for visualization into a single software architecture, under the assumption of an iterative workflow of user actions and result inspection. In this work, we propose a unified architecture intended for use by researchers and developers. While this architecture is designed for an $N$-dimensional space, the present study implements and evaluates it in 4D.
The contribution of this work is to present an architectural configuration that integrates the above processing components as a single software architecture and to evaluate the 4D implementation experimentally.
In this paper, interactivity is defined as a workflow that presupposes repeated cycles of user actions and result inspection, in contrast to offline batch processing in which input conditions are fixed in advance and executed in a single run. Hereafter, the term “interactive,” including its adjectival and adverbial usages, refers to usage scenarios that satisfy or require this notion of interactivity.
@@ 2 Related Work
Prior work on visualization and simulation in high-dimensional spaces is often reported in terms of individual elemental techniques.
For convex hull computation applicable to high dimensions, the pivoting method [2] and Quickhull [3] are known as representative approaches. In implementations that ensure robustness against inputs with numerical errors or degenerate configurations, these algorithms can incur increased computational cost due to exact predicates and additional handling.
Convex hull computation is part of the fundamental problem set in high-dimensional computational geometry, and textbooks systematically organize problems such as convex hulls, halfspace intersection, and arrangements [18].
As existing methods for high-dimensional visualization, there are approaches based on projection to lower-dimensional spaces. Cook et al. provide an overview of methods and design issues for exploring and presenting high-dimensional information via 2D projections [6]. Dimensional stacking has also been surveyed as a class of visualization techniques that assigns each dimension to a nested 2D layout [12].
These works mainly survey visualization techniques. In contrast, on the implementation side, architectures and frameworks for high-dimensional visualization have also been proposed. For mesh-based approaches, a GPU-based 4D visualization architecture has been proposed [5] and discussed from the viewpoint of rendering and visualization of 4D shapes. As a software framework for computations on convex polyhedra, “polymake” has been reported [10]. It provides visualization, exemplified by Schlegel diagrams, together with the underlying algorithmic processing. However, these prior examples differ in scope from the unified environment targeted in this paper, which aims to integrate not only visualization but also other processing in an interactive manner.
For Boolean operations on 3D meshes, methods based on intersection resolution and spatial partitioning have been reported; Zhou et al. present a unified framework that handles union, intersection, and difference based on cell decomposition and extraction using winding number vectors [26]. In the 3D setting, implementation examples of interactive Boolean operations have also been reported [4]. CGAL adopts a design policy that emphasizes geometric robustness [8]. Under this policy, CGAL and related libraries provide set operations in 2D and 3D, exemplified by Nef polyhedra, whereas implementation reports of Boolean operations on explicit boundary meshes in 4D and higher dimensions are limited.
For set operations in 4D and higher dimensions, libraries for high-dimensional computation based on polyhedral representations have been proposed [24]. However, these assume representations based on halfspace constraints or cell decompositions, rather than explicit boundary meshes. They also typically assume offline computation with exact arithmetic, and implementation reports aimed at interactive operating environments remain limited.
In physical simulation, non-rigid body simulation based on XPBD has been reported in 3D space [15], and there are also studies that handle rigid bodies in 4D and higher dimensions [22]. However, reports whose primary goal is to visualize non-rigid objects in 4D and higher dimensions and handle them interactively remain limited.
@@ 3 Proposed Method
3.1 Overview of the Overall Architecture
Figure 1: An overview of the proposed architecture. The left side shows the static shape definition workflow, and the right side shows the dynamic visualization pipeline.
This section describes the technical details of the proposed architecture. We first provide an overview of the overall architecture, and then describe in detail the algorithms that constitute each function and their implementations.
As shown in Fig. 1, the overall structure and data flow of the architecture consist of multiple components: convex hull mesh generation, Boolean operations, coordinate transformations, and hyperplane slicing. The convex hull mesh generation component employs a Quickhull-based algorithm to construct a list of facet indices that define the topology of an $N$-dimensional object from the input point set data. The Boolean operation component dynamically modifies the topology and geometry of the mesh based on user interactions. The coordinate transformation component converts vertex coordinates into a camera-referenced coordinate system by applying a view transformation based on the viewpoint, after incorporating orientation transformations updated through user operations. The hyperplane slicing component applies hyperplane slicing to the output mesh from the coordinate transformation component and generates and renders 3D cross-sections within the update loop.
In this study, geometric algebra is adopted as the mathematical foundation for uniformly describing geometric entities such as $N$-dimensional vectors, planes, and rotations. Introductory explanations of geometric algebra and its use in computer science and computer graphics can be found in [23], systematic treatments in [7], and survey-style overviews in [14].
@@ 3.2 Design Philosophy and Architectural Overview
The design policy of this architecture emphasizes three principles: algorithmic transparency, minimization of dependence on external libraries, and separation of topology and geometry. This policy is based on a design decision that prioritizes simplification of description and implementation over performance optimization, so that researchers and developers can trace the structure of the algorithms.
Based on this design policy, all computational geometry logic is unified under a CPU-based implementation written in standard C#. While GPU-based parallelization can deliver high performance for specific tasks, it introduces issues of code complexity and hardware dependency. The CPU-based C# implementation is intended to enhance platform independence and algorithmic transparency. In addition, this paper adopts an approximate implementation that tolerates numerical errors and does not guarantee numerical robustness arising from floating-point arithmetic, which is positioned as a subject for future work. The 4D implementation in this study was developed in C# on the Unity engine and is released as open-source software [1].
In this architecture, a design is adopted in which the vertex index sequences that define facets (topology) and the vertex coordinate arrays (geometry) are managed independently. This design assumes scenarios in which time-varying geometry is provided as input for a fixed topology. Details are described in Section 3.8.
3.3 N-Dimensional Convex Hull Mesh Generation
@@ 3.3.1 Comparison with Existing Methods
This component is centered on a self-contained C# implementation of the classical $N$-dimensional Quickhull algorithm [3]. A standard implementation would perform an optimization called candidate point pruning to maximize computational efficiency. In this strategy, each facet maintains its own list of candidate points, and points that have been absorbed into the interior of the convex hull are excluded from further computation; this is effective for volume-like point sets containing many interior points. However, in this architecture, the primary use cases are basic convex hull mesh construction for initial shapes and dynamic simplex tessellation (described in Section 3.3.3), and it often handles small surface-like point sets with only a few dozen vertices and few interior points. In this context, the performance gains from pruning are limited, whereas increased implementation complexity entails a risk of reduced code maintainability.
Therefore, this architecture deliberately removes the pruning mechanism in light of this trade-off. As shown in Fig. 2, it linearly searches a single global set of candidate points to find the farthest point, and removes from the candidate set only those points that are used as convex hull vertices. In this paper, we refer to this approach as “Direct Quickhull.”
@@ 3.3.2 Calculating N-Dimensional Normals and Visible Facets
In Direct Quickhull, facet normal vectors are computed iteratively, and this computation can become a dominant factor in the overall processing time. One possible approach for computing normals is to use null-space computation; however, this typically requires singular value decomposition (SVD) or similar methods, which can impose a substantial computational burden in the iterative calculations assumed for interactive use. To address this issue, this architecture introduces the concepts of the wedge product and the dual from geometric Algebra. First, in an $N$-dimensional space $\mathbb{R}^{N}$, we define a facet from its $N$ vertices $\{\mathbf{p}_{0},\mathbf{p}_{1},\dots,\mathbf{p}_{N-1}\}$ and construct the (N1)(N-1) linearly independent vectors that span the facet.
@@
Next, we calculate the wedge product of these vectors.
@@
In geometric Algebra, this (N1)(N-1)-blade $\mathbf{B}$ is a mathematical entity that represents the oriented hyperplane spanned by the vectors. The operation of taking the Hodge dual of this blade, $\mathbf{B}^{*}$, is equivalent to finding the orthogonal complement of that hyperplane. That is, for an (N1)(N-1)-dimensional hyperplane, its orthogonal complement is a 1-dimensional line, and its direction vector becomes the normal vector $\mathbf{n}$ to the hyperplane.
@@
Next, it is necessary to correctly determine the orientation of the facet’s normal vector. In the 3D case, the normal vectors of facets constituting the surface of a convex hull were obtained with correct orientation by computing the ordering of vertices, exploiting the right-handed property of the cross product of 3D vectors. However, the cross product is unique to 3D space. In higher dimensions, there are no geometric concepts analogous to the winding order that intuitively define a face’s orientation.
Therefore, in this architecture, this issue is addressed by exploiting the geometric property that facet normal vectors of a convex hull always point outward. The orientation is determined by computing the dot product between the facet normal vector and a vector directed from a point on the facet toward the centroid of the initial simplex constructed at the beginning of the algorithm. If the sign of the dot product is negative, the normal vector is judged to be outward-facing; if the sign is positive, the normal vector is flipped to obtain the correct outward orientation.
Next, we use the facet normal vectors to determine the visible facets, which is a necessary step in the process of expanding the convex hull.
@@
3.3.3 Simplex Filling in Convex Hull Construction
In this architecture, Direct Quickhull is used for simplicial decomposition of the convex hull. In the incremental updates of Direct Quickhull, the newly added region is represented as a set of $N$-dimensional simplices generated from the new vertex $\mathbf{P}_{\text{new}}$ and the visible facets. Exploiting this property, the architecture appends the simplices generated at each update step to a simplex list and retains them, thereby obtaining the boundary facet set and the simplex list upon completion of convex hull construction.
This method is used for two purposes. First, it is used for the simplicial decomposition of a convex polytope produced by hyperplane clipping of facets in Boolean operations. Second, it is intended for determining the cross-sectional structure of facets arising in hyperplane slicing in spaces of five or more dimensions; details are described in Section 3.6.3.
@@ 3.4.1 Broad Phase: Candidate Pruning via Spatial Partitioning
In $N$-dimensional Boolean operations, let the numbers of facets in the two meshes be $N_{\text{A}}$ and $N_{\text{B}}$, respectively. A brute-force intersection test over all facet pairs then has a computational complexity of $O(N_{\text{A}}N_{\text{B}})$. To avoid brute-force intersection testing, this architecture adopts a spatial partitioning grid as a broad-phase stage and extracts facet pairs that are potential intersection candidates. We compute the axis-aligned bounding box (AABB) for each facet and register the facet with the grid cells that its AABB occupies. In the subsequent narrow-phase stage, intersection tests are performed on facet pairs that share the same cell. Hierarchical methods such as BVHs and octrees could also be considered as alternatives; however, this paper prioritizes implementation simplicity and adopts a uniform grid approach.
@@ 3.4.2 Narrow Phase: Facet-Facet Intersection Test
In the narrow phase, we perform precise geometric intersection tests on the candidate pairs pruned during the broad phase. The principle of this algorithm is a generalization to $N$-dimensions of the method proposed by Tomas Möller et al. for fast triangle-triangle intersection tests in 3D [17]. Its essence lies in decomposing the intersection test between two facets, $\mathcal{F}_{\text{A}}$ and $\mathcal{F}_{\text{B}}$, into two symmetric sub-problems. First, we compute the geometric intersection region, $\mathcal{I}_{\text{B}}$, between the (N1)(N-1)-dimensional hyperplane spanned by $\mathcal{F}_{\text{A}}$ and the facet $\mathcal{F}_{\text{B}}$. Second, we perform the same operation with the roles swapped, computing the intersection region $\mathcal{I}_{\text{A}}$ between the hyperplane spanned by $\mathcal{I}_{\text{B}}$ and the facet $\mathcal{F}_{\text{A}}$. Both of these intersection regions are (N2)(N-2)-dimensional convex polytopes. Finally, we determine if an intersection between the facets exists by testing whether $\mathcal{I}_{\text{A}}$ and $\mathcal{I}_{\text{B}}$ overlap within their common (N2)(N-2)-dimensional space.
If an intersection is found to exist, this information is utilized in the subsequent tessellation stage. Specifically, we register this intersection pair in a dictionary where one facet is the key and the other is the value. This registration is performed for all intersecting pairs, and the cutting relationships are represented as a dictionary, including cases in which a single facet has multiple intersection counterparts.
In the 4D implementation, the Separating Axis Theorem (SAT) is adopted to determine overlap between intersection regions. However, in dimensions of five or higher, the number of vertices of the (N2)(N-2)-dimensional polytopes representing intersection regions can increase, which in turn increases the number of candidate separating axes enumerated by SAT and may lead to higher computational cost. To address this issue, for applications in five or more dimensions, an alternative candidate is intersection testing based on the GJK algorithm.